A computer monitor has a height of 12 inches with a diagonal measurement of 20 inches. How many inches wide is the rectangular screen ?
Answer by unlockmath(1688)

Hello,
Use this formula to solve.
A^2 + B^2 = C^2
Plug in the numbers:
12^2 + B^2 = 20^2
Rewrite:
144 + B^2 = 400
Subtract 144 and sq rt and we get
16 inches
Make sense?
RJ
www.math-unlock.com